# Text2PDE: Latent Diffusion Models for Accessible Physics Simulation

### v1

@@ 3.3 Conditioning Mechanisms
To sample from the conditional distribution $p(\mathbf{u}|\mathbf{u}^{0},B[\mathbf{u}])$, we add the conditional information to the denoising model $\epsilon_{\theta}(\mathbf{z_{n}},n,\mathbf{u}^{0},B[\mathbf{u}])$. This adds the additional consideration of how to represent initial and boundary conditions in latent space. For current PDE benchmarks and in many practical applications, the boundary conditions do not change between solutions; as such, we use $\mathbf{u}(0,\mathbf{x}_{m})$ as a proxy for both $\mathbf{u}^{0}$ and $B[\mathbf{u}]$. As such, conditioning information can be added to the denoising backbone by defining a domain-specific encoder $\tau_{\theta}$ that maps the first frame of the solution to a conditioning sequence $\mathbf{c}=\tau_{\theta}(\mathbf{u}(0,\mathbf{x}_{m}))\in\mathbb{R}^{N_{c}\times d_{c}}$, where $N_{c}$ is the sequence length, and $d_{c}$ is the conditioning dimension. Conveniently, the proposed encoder $\mathcal{E}$ is agnostic to the temporal and spatial discretization of PDEs; therefore, the same architecture can be used for $\tau_{\theta}$ by flattening the downsampled latent dimension $(T_{l}/f\times M_{l}/f=N_{c})$.
@@
Another interesting conditioning domain is natural language. Although imprecise, humans describe and understand physics through language; for example, the speed of a projectile (fast/slow), or the mass of an object (heavy/light). In principle, the initial and boundary conditions of a physics simulation could also be described by text, for example: a large plume of smoke (initial condition) entering a small box (boundary condition). An advantage of this modality is that the physical simulation can be easily prescribed; rather than needing to know the discretization scheme and the physical values at every mesh point, only the important behavior driving the physical phenomena need to be described. This has the additional benefit of removing redundancies in PDE data, as not all physical values are important, as well as improving the interpretability and usability of physics simulators. However, a clear drawback is the underdetermined nature of the PDE problem: for a given prompt describing the initial and boundary conditions, there may exist many plausible solutions. Furthermore, this approach introduces the nontrivial problem of captioning physics simulations; we develop additional methods to achieve this and interested readers are directed to Appendix D for details on PDE captioning.
@@
Model Params Tflops L1 Loss
GINO 72M 0.73 0.0625
MGN 101M 32.16 0.0712
OFormer 131M 17.34 0.0750
LDMS-FF 198M 0.81 0.0407
LDMM-FF 667M 1.16 0.0385
LDMS-Text 313M 0.83 0.0470
LDMM-Text 804M 1.20 0.0404
Table 1: Cylinder Flow. Left: Model parameters, flops, and validation losses across different benchmarks. Right: Sample rollouts for first frame and text conditioned models compared to the reference solution.
@@ 4 Experiments
We explore the proposed latent diffusion model (LDM) for two 2D PDE datasets: flows around a cylinder (Pfaff et al., 2021) and buoyancy-driven smoke flows (Gupta & Brandstetter, 2022). Both datasets model the Navier-Stokes equations, although with different regimes, boundary conditions, solution domains, and physical variables. In particular, the cylinder flow dataset is discretized on a mesh to smoothly model the cylindrical geometry and its wake; however, the smoke buoyancy dataset is uniformly discretized in order to benchmark the proposed model against regular-grid neural surrogates, which are more common. Additional dataset details can be found in Appendix E. For each model, we report its parameter count and L1 loss. Although commonly reported, parameter count is a poor measure of training cost and model complexity (Peebles & Xie, 2023); therefore, we additionally report the floating point operations per second (flops) needed for a single forward pass during training with a batch size of 1, calculated using DeepSpeed (Rasley et al., 2020). Additional details on training and model hyperparameters can be found in Appendix F.
@@ 4.1 Cylinder Flow
Following Pfaff et al. (2021), we use 1000 samples for training and 100 samples for validation. Each sample contains approximately 2000 mesh points and is downsampled to evolve over 25 timesteps, which is compressed to a latent size of $16\times 16\times 16$. Each sample is varied in the position and radius of the cylinder, as well as in the inlet velocity of the fluid. This results in samples with Reynolds numbers of around $10^{2}\sim 10^{3}$, which for this problem span the laminar and transitional regimes (Stringer et al., 2014). Lastly, the physical variables are velocity and pressure.
Interestingly, for this dataset, a text prompt can describe a solution well. The problem setup assumes that all parameters are fixed except for the cylinder radius, cylinder position, and inlet velocity, which can be conveniently inserted into a text prompt. In particular, to caption this dataset, we extract these values and procedurally generate prompts based on a template; for interested readers, we include further details and ablation studies in Appendix D.1.
@@ Results
We consider benchmarking against GINO (Li et al., 2023c), MeshGraphNet (MGN) (Pfaff et al., 2021), and OFormer (Li et al., 2023a), which represent a variety of operator-, graph-, and attention-based neural surrogates. Since these are autoregressive models, we provide the initial frame of the solution and autoregressively predict the next 24 frames. We compare this to our latent diffusion model conditioned on both the first frame (-FF) and a text prompt (-Text), which are trained to generate a full solution trajectory at once. Furthermore, we consider two model sizes: Small (S) and Medium (M); the results are shown in Table 1.
We observe that the latent diffusion model can outperform current models on this benchmark. Notably the inclusion of flops allows for additional insight into model efficiency. We reproduce results showing that GINO models are very efficient (Li et al., 2023c), yet LDMs can still match this efficiency by using a latent space and diffusing an entire rollout at once. Furthermore, our model is able to achieve lower errors than all baselines and more efficiently than graph- or attention-based neural surrogates. Lastly, conditioning on the first frame outperforms text conditioning, presumably because more information is provided in the initial velocity and pressure fields. However, text still remains an accurate conditioning modality; moreover, the latent diffusion model shows good scaling behavior, with increased performance at larger model sizes.
Model Params Tflops L1 Loss
FNO 510M 0.85 0.2134
Unet 580M 21.48 0.1287
Dil-Resnet 33.2M 58.75 0.1964
ACDM 404M 51.23 0.2063
LDMS-FF 243M 1.41 0.1540
LDMM-FF 725M 1.68 0.1386
LDML-FF 1.55B 2.19 0.1178
LDMS-Text 334M 1.39 0.1738
LDMM-Text 825M 1.65 0.1320
LDML-Text 2.69B 2.53 0.1240
@@ 4.2 Buoyancy-Driven Flow
We additionally evaluate our model on a regular-grid benchmark of conditional buoyancy-driven smoke rising in a square domain (Gupta & Brandstetter, 2022). We use 2496 training samples and 608 validation samples, with varying initial conditions and buoyancy factors, and at a spatial resolution of $128\times 128$ and 48 timesteps, which is compressed to a latent dimension of $6\times 16\times 16$. The smoke density is modeled by a passive transport equation, and as a result, the physical variables of interest are velocity in 2 dimensions and density. Since the flow is more constrained by its boundaries, this benchmark contains smaller-scale features and more complex mixing behaviors.
For the smoke buoyancy problem, it is challenging to describe an initial plume of smoke with text. In particular, plumes may exhibit different shapes, locations, and sizes, in addition to the initial velocity field being even less visually distinct. Nevertheless, we investigate captioning the initial condition using a multimodal LLM. We design a prompt and provide the density field as an image; additionally, we outline individual smoke plumes in the initial density field using a Canny edge detector (Canny, 1986) and provide this segmented image in the LLM prompt as well. For additional details and examples, we refer readers to Appendix D.2.
@@ Results
We consider benchmarking against the popular FNO, Unet, and Dilated Resnet architectures (Li et al., 2021; Ronneberger et al., 2015; Gupta & Brandstetter, 2022; Stachenfeld et al., 2022). Additionally, we benchmark against an auto-regressive diffusion model (ACDM) (Kohl et al., 2024), to evaluate the need for a latent space and full spatio-temporal sample generation. To ensure a fair comparison, we only provide these baselines with the initial solution and autoregressively predict the next 47 frames. We compare this to our proposed model, conditioned on the first frame (-FF) as well as the text prompt (-Text), and in different model sizes: Small (S), Medium (M), and Large (L). The results are given in Table 2. Some higher-frequency details are lost in the LDM prediction; this could be due to working in a latent space that is 512 times smaller than the original dimension.
@@
Due to the underdetermined nature of text conditioning, we propose an additional evaluation metric. This is motivated by the observation that text-conditioned diffusion models often generate a variety of initial conditions that semantically satisfy a given prompt, yet this can affect the temporal rollout significantly. Although the sampled solution may not match a specific solution in the validation set, it still fulfills the semantic goal of text conditioning. Therefore, use a numerical solver to evolve initial conditions taken from a sampled solution to produce the ground truth. This would evaluate the physical consistency of generated solutions without restricting the diversity of the LDM simulator. We implement this by sampling a solution for each text prompt in the validation set, then using the sampled initial condition to produce a ground-truth using PhiFlow (Holl & Thuerey, 2024). The results of this evaluation are also reported in Table 2, with additional results in Appendix A.
@@
We observe that LDM models conditioned on the first frame of a PDE solution can outperform current models and are an efficient class of neural surrogates. Furthermore, text-conditioned diffusion models are able to model a set of phenomena that both semantically satisfy the prompt and are physically consistent, although the samples may not match a specific sample in the validation set. An interesting observation is that many text-conditioned samples are not contained in the training or validation sets, indicating that the generative model has learned the underlying physics to generate entirely new samples based on a text prompt. Furthermore, we note that for this specific benchmark, initial conditions are generated from a Gaussian distribution, which is inherently challenging to describe with a text prompt. Real-world problems often contain more structure that natural language can better describe, such as in the cylinder flow problem. Lastly, the LDM model displays good scaling behavior, and even at large parameter counts, its training remains efficient.
@@ 4.3 Discussion
We show that latent diffusion models can be a viable approach to generating spatio-temporal physics simulations. Specifically, by sampling spatio-temporal noise and leveraging the ability of diffusion models to approximate complex distributions, an entire trajectory can be generated from an initial condition. This circumvents the traditional paradigm whereby solvers need to evolve solutions based on a current timestep, and aids in increasing accuracy by preventing error accumulation. We demonstrate this by plotting the time-dependent prediction error for various models in Figure 3. Additionally, the use of a latent space allows this computation to be performed efficiently and with unstructured data. Diffusion models generally assume image-like spatially invariant data, which is usually detrimental in physics applications; the introduction of a mesh autoencoder allows latents to be denoised and decoded where solutions need to be refined (i.e. wake regions) and onto arbitrary geometries. Lastly, the scaling behavior of the proposed model is promising and reflects similar observations of scaling diffusion and attention-based models in other domains (Peebles & Xie, 2023; Kaplan et al., 2020). The benchmarked datasets are relatively small (<5000) samples; with a larger, more diverse dataset, this scaling ability could unlock unified physics surrogates across a wide variety of phenomena.
Additionally, the introduction of text-conditioned physics simulators is an interesting and novel research direction. In constrained PDE formulations, text is a compact and accurate conditioning modality, nearly matching first-frame conditioning in cylinder flow problems despite conditioning on around 1000×1000$\times$1000 × less data (${\sim}1$Mb vs. ${\sim}1$Kb), and with the added advantage of being more interpretable and usable. In more practical scenarios where the input space is less defined, the diversity and ease of text-conditioned physics simulators could be a useful tool for engineers and designers to quickly test different ideas without needing to define geometries, discretizations, or generate an initial solution. After deciding on a suitable design, the same architecture can then be used with first-frame conditioning to more accurately model the resulting physical phenomena.
@@ 5.1 Video Diffusion Models
Many prior works have considered using diffusion models to generate videos. Initial works have focused on adapting image diffusion models to the video domain, proposing many mechanisms to ensure temporal consistency or upsample spatio-temporal resolutions (Singer et al., 2022; Wu et al., 2023; Blattmann et al., 2023b). Concurrently, many approaches have investigated directly generating videos by modifying the diffusion backbone with a 3D-Unet (Ho et al., 2022b), with subsequent works improving the resolution, quality, and consistency of generated videos (Ho et al., 2022a; Bar-Tal et al., 2024). Within these two paradigms, an abundance of research has examined the scaling of video diffusion models (Blattmann et al., 2023a), improved conditioning methods (Girdhar et al., 2023), training-free adaptation (Khachatryan et al., 2023; Zhang et al., 2023), and video editing (Esser et al., 2023).

### v2

@@ 3.3 Conditioning Mechanisms
To sample from the conditional distribution $p(\mathbf{u}|\mathbf{u}^{0},B[\mathbf{u}])$, we add the conditional information to the denoising model $\epsilon_{\theta}(\mathbf{z_{n}},n,\mathbf{u}^{0},B[\mathbf{u}])$. For current PDE benchmarks and in many practical applications, the boundary conditions do not change between solutions; as such, we use $\mathbf{u}(0,\mathbf{x}_{m})$ as a proxy for both $\mathbf{u}^{0}$ and $B[\mathbf{u}]$. As such, conditioning information can be added to the denoising backbone by defining a domain-specific encoder $\tau_{\theta}$ that maps the first frame of the solution to a conditioning sequence $\mathbf{c}=\tau_{\theta}(\mathbf{u}(0,\mathbf{x}_{m}))\in\mathbb{R}^{N_{c}\times d_{c}}$, where $N_{c}$ is the sequence length, and $d_{c}$ is the conditioning dimension. Conveniently, the proposed encoder $\mathcal{E}$ is agnostic to the temporal and spatial discretization of PDEs; therefore, the same architecture can be used for $\tau_{\theta}$ by flattening the downsampled latent dimension $(T_{l}/f\times M_{l}/f=N_{c})$.
@@
Another interesting conditioning domain is natural language. Although imprecise, humans describe and understand physics through language, and in principle, the initial and boundary conditions of a physics simulation could also be described by text. An advantage of this modality is that it is interpretable and compact; instead of needing to specify the discretization scheme and the physical values at every mesh point, only the important behavior driving the physical phenomena need to be described. However, a clear drawback is the underdetermined nature of the PDE problem: for a given prompt describing the initial and boundary conditions, there may exist many plausible solutions. Furthermore, this approach introduces the problem of captioning physics simulations; we develop additional methods to achieve this and interested readers are directed to Appendix D for details on PDE captioning.
@@
Model Params Tflops L2 Loss
GINO 72M 0.73 0.2445
MGN 101M 32.16 0.2617
OFormer 131M 17.34 0.3386
LDMS-FF 198M 0.81 0.1522
LDMM-FF 667M 1.16 0.1309
LDMS-Text 313M 0.83 0.1796
LDMM-Text 804M 1.20 0.1476
Table 1: Cylinder Flow. Left: Model parameters, flops, and validation losses across different baselines. Right: Sample rollouts for first frame and text conditioned models compared to the reference solution.
@@ 4 Experiments
We explore the proposed latent diffusion model (LDM) for three PDE datasets: 2D flows around a cylinder (Pfaff et al., 2021), 2D buoyancy-driven smoke flows (Gupta & Brandstetter, 2022), and 3D turbulence around geometric objects (Lienen et al., 2024). The cylinder flow dataset is discretized on a mesh to smoothly model the cylindrical geometry and its wake; however, the smoke buoyancy and 3D turbulence datasets are uniformly discretized. Additional dataset details can be found in Appendix E. For each model, we report its parameter count and relative L2 loss. Since parameter count can be a poor measure of training cost and model complexity (Peebles & Xie, 2023), we additionally report the floating point operations per second (flops) needed for a single forward pass during training with a batch size of 1, calculated using DeepSpeed (Rasley et al., 2020). Additional details on training and model hyperparameters can be found in Appendix F.
@@ 4.1 Cylinder Flow
Following Pfaff et al. (2021), we use 1000 samples for training and 100 samples for validation. Each sample contains approximately 2000 mesh points and is downsampled to evolve over 25 timesteps, which is compressed to a latent size of $16\times 16\times 16$. Each sample is varied in the position and radius of the cylinder, as well as in the inlet velocity of the fluid. This results in samples with Reynolds numbers of 100-1500 and includes Karman vortex streets (Stringer et al., 2014b). In addition, the physical variables are velocity and pressure. Lastly, to caption this dataset, we extract the cylinder position, radius, and Reynolds number and procedurally generate prompts based on a template; we include further details and ablation studies in Appendix D.1.
@@ Results
We consider benchmarking against GINO (Li et al., 2023c), MeshGraphNet (MGN) (Pfaff et al., 2021), and OFormer (Li et al., 2023a), which represent a variety of operator-, graph-, and attention-based neural solvers. Since these are autoregressive models, we provide the initial frame of the solution and autoregressively predict the next 24 frames. We compare this to our latent diffusion model conditioned on both the first frame (-FF) and a text prompt (-Text), which are trained to generate a full solution trajectory at once. Furthermore, we consider two model sizes: Small (S) and Medium (M); the results are shown in Table 1 with relative L2 loss.
We observe that the latent diffusion model can outperform current models on this benchmark and pressure errors are around the same as velocity errors. Notably the inclusion of flops allows for additional insight into model efficiency. We reproduce results showing that GINO models are very efficient (Li et al., 2023c), yet LDMs can still match this by using a latent space and diffusing an entire rollout at once. Furthermore, our model can achieve lower errors than baselines and more efficiently than graph- or attention-based neural surrogates. Lastly, conditioning on the first frame outperforms text conditioning, presumably because more information is provided in the initial velocity and pressure fields. Text remains an accurate conditioning modality; moreover, the latent diffusion model shows good scaling behavior, with increased performance at larger model sizes.
Model Params Tflops L2 Loss
FNO 510M 0.85 0.5126
Unet 580M 21.48 0.3050
Dil-Resnet 33.2M 58.75 0.4466
ACDM 404M 51.23 0.4766
LDMS-FF 243M 1.41 0.3459
LDMM-FF 725M 1.68 0.3177
LDML-FF 1.55B 2.19 0.2728
LDMS-Text 334M 1.39 0.4290
LDMM-Text 825M 1.65 0.3158
LDML-Text 2.69B 2.53 0.2944
@@ 4.2 Buoyancy-Driven Flow
We additionally evaluate our model on a regular-grid benchmark of conditional buoyancy-driven smoke rising in a square domain (Gupta & Brandstetter, 2022). We use 2496 training samples and 608 validation samples, with varying initial conditions and buoyancy factors, and at a spatial resolution of $128\times 128$ and 48 timesteps, which is compressed to a latent dimension of $6\times 16\times 16$. The smoke density is modeled by a passive transport equation, and as a result, the physical variables are velocity and density. Lastly, for the smoke buoyancy problem, it is more challenging to describe an initial condition with text. In particular, plumes may exhibit different shapes, locations, and sizes, in addition to the initial velocity field being even less visually distinct. Nevertheless, we investigate captioning the initial condition using a multimodal LLM. We design a prompt and provide the density field as an image; additionally, we outline individual smoke plumes in the initial density field using a Canny edge detector (Canny, 1986) and provide this segmented image in the LLM prompt as well. For additional details and examples, we refer readers to Appendix D.2.
@@ Results
We consider benchmarking against the FNO, Unet, and Dilated Resnet architectures (Li et al., 2021; Ronneberger et al., 2015; Gupta & Brandstetter, 2022; Stachenfeld et al., 2022). Additionally, we benchmark against an autoregressive diffusion model (ACDM) (Kohl et al., 2024), to evaluate the need for a latent space and full spatio-temporal sample generation. To ensure a fair comparison, we provide these baselines with the initial solution and autoregressively predict the next 47 frames. We compare this to our LDM model, conditioned on the first frame (-FF) or the text prompt (-Text), and in different model sizes: Small (S), Medium (M), and Large (L), in Table 2.
@@
Due to the underdetermined nature of text conditioning, we propose an additional evaluation metric. This is motivated by the observation that text-conditioned diffusion models often generate a variety of initial conditions that semantically satisfy a given prompt, yet this can affect the temporal rollout significantly. Although the sampled solution may not match a specific solution in the validation set, it still fulfills the semantic goal of text conditioning. Therefore, we sample a solution for each text prompt in the validation set, then use the sampled initial condition to produce a ground-truth using PhiFlow (Holl & Thuerey, 2024), which would evaluate the physical consistency of generated solutions without restricting the diversity of the LDM simulator. The losses using this re-solved trajectory are reported in Table 2, with additional visualizations in Appendix A.
@@
We observe that LDM models conditioned on the first frame of a PDE solution can outperform current models and are an efficient class of neural solvers. Additionally, density errors are around 1.5 times the velocity errors. Furthermore, text-conditioned diffusion models are able to model a set of phenomena that both semantically satisfy the prompt and are physically consistent, although the samples may not match the validation set. In addition, many text-conditioned samples are not contained in the training or validation sets, indicating that the generative model has learned the underlying physics to generate new samples based on a text prompt. Furthermore, for this benchmark, the initial conditions are generated from a Gaussian distribution, which is challenging to describe with a prompt and real-world problems often contain more structure that natural language can better describe. Lastly, the LDM model displays good scaling behavior, with efficient training at large parameter counts.
Model Params Tflops L2 Loss $D_{TKE}$
FNO 1.02B 1.62 0.862 6.524
FactFormer 41.4M 53.8 0.795 6.022
Dil-Resnet 24.8M 249 0.707 6.153
LDM-FF 2.72B 9.78 0.602 5.630
LDM-Text 2.73B 9.43 0.693 5.653
Table 3: 3D Turbulence. Left: Model parameters, flops, validation losses and log-TKE distance. Right: Sample rollouts for first frame and text conditioned models compared to the reference solution.
4.3 3D Turbulence
Dataset
To investigate scaling behavior to larger, more complex systems, we consider evaluating the LDM framework on a 3D turbulence problem of air flows over various geometric objects such as squares, elbows, U-shapes, etc. (Lienen et al., 2024). We use 36 training samples and 9 validation samples, downsampled to a resolution of $96\times 24\times 24$ and 1000 timesteps; however, we set the prediction horizon to 48 timesteps during evaluation. The average Reynolds number is $52e$ 5 and the physical variables are velocity and pressure. Additionally, the simulations are manually captioned based on the geometry of the initial condition; we include additional details in Appendix D.3.
Results
We consider benchmarking against FNO3D (Li et al., 2021) as well as FactFormer (Li et al., 2023b), a scalable attention-based neural solver, and Dil-Resnet (Stachenfeld et al., 2022), an accurate yet computationally expensive CNN solver. In addition to relative L2 loss, we report the L2 distance between the log turbulent kinetic energy $D_{TKE}$ between predicted and true samples, since chaotic and small-scale fluctuations in turbulence make it unlikely for similar samples to be close under Euclidean distance when evolved over time (Lienen et al., 2024). It is currently challenging for any neural surrogate to accurately model 3D turbulence, however the proposed model is able to do so more accurately than baselines and remains efficient. Although first-frame conditioning is more accurate, text conditioning is still an accurate conditioning modality. Lastly, there are significant smoothing effects which seems to be the cost for temporal stability and efficient training in 3D.
@@ 4.4 Discussion
We show that latent diffusion models can be a viable approach to generating spatio-temporal physics simulations. Specifically, by sampling spatio-temporal noise and leveraging the ability of diffusion models to approximate complex distributions, an entire trajectory can be generated from an initial condition. This circumvents the traditional paradigm whereby solvers need to evolve solutions based on a current timestep, and aids in increasing accuracy by preventing error accumulation. We demonstrate this by plotting the time-dependent prediction error for various models in Figure 3. Additionally, the use of a latent space allows this computation to be performed efficiently and with unstructured data, and the introduction of a mesh autoencoder allows latents to be decoded where solutions need to be refined (i.e. wake regions) and onto arbitrary geometries. Lastly, the scaling behavior of the proposed model is promising and reflects similar observations of scaling diffusion and attention-based models in other domains (Peebles & Xie, 2023; Kaplan et al., 2020). The benchmarked datasets are relatively small (<5000) samples; with a larger, more diverse dataset, this scaling ability could unlock unified physics surrogates across a wide variety of phenomena.
Additionally, the introduction of text-conditioned physics simulators is an interesting and novel research direction. In constrained PDE formulations, text is a compact and accurate conditioning modality, nearly matching first-frame conditioning in cylinder flow problems despite conditioning on around 1000×1000$\times$1000 × less data (${\sim}1$Mb vs. ${\sim}1$Kb), and with the added advantage of being more interpretable and usable. The diversity and ease of text-conditioned physics simulators could be a useful tool for engineers and designers to quickly test different ideas without needing to define geometries, discretizations, or generate an initial solution. After deciding on a design, the same architecture can then be used with first-frame conditioning to precisely model the resulting physical phenomena.
@@ 5.1 Video Diffusion Models
Many prior works have considered using diffusion models to generate videos. Initial works have focused on adapting image diffusion models to the video domain, proposing mechanisms to ensure temporal consistency or upsample spatio-temporal resolutions (Singer et al., 2022; Wu et al., 2023; Blattmann et al., 2023b). Many approaches have investigated directly generating videos by modifying the diffusion backbone (Ho et al., 2022b), with subsequent works improving the resolution, quality, and consistency of generated videos (Ho et al., 2022a; Bar-Tal et al., 2024). Within these two paradigms, an abundance of research has examined the scaling of video diffusion models (Blattmann et al., 2023a), improved conditioning methods (Girdhar et al., 2023), training-free adaptation (Khachatryan et al., 2023; Zhang et al., 2023), and video editing (Esser et al., 2023).